OpenStudy (anonymous):

please help. What is the slope of a line perpendicular to the line 12x - 4y = 7? A. -7/4 B. 7/4 C. -1/3 D. 1/3

OpenStudy (anonymous):

@primeralph

OpenStudy (anonymous):

well what is the slope of the line given?

OpenStudy (anonymous):

is it A ? - 7/4

OpenStudy (anonymous):

no, it's not A

OpenStudy (nurali):

To find the slpoe of a line we use the following formula: Slope of the line= -x/y Now come to the question Given line is 12x - 4y = 7 Let "m" be the slope of the given line, then m = -12 / -4 m = 3 Now there is another line which is perpendicular to the given line, so let m' be the slope of the perpendicular line. By defination, m x m' = -1 3 x m' = -1 m' = -1 / 3

OpenStudy (anonymous):

would have never expected it to be -1/3 ... but thanks.

OpenStudy (anonymous):

well if you know that the slope of the given line is 3... then -1/3 would seem pretty expected right?

OpenStudy (anonymous):

how am i suppose to know tje slope of the line?

OpenStudy (anonymous):

i would have thought the slope was 7